# Long-Sequence Recommendation Models Need Decoupled Embeddings

### v1

@@ 4.1 Setup
We use the publicly available Taobao (Zhu et al., 2018; 2019; Zhuo et al., 2020) and Tmall (Tianchi, 2018) datasets, which provide users’ behavior data over specific time periods on their platforms. Each dataset includes the items users clicked, represented by item IDs and their corresponding category IDs. Thus, a user’s history is modeled as a sequence of item and category IDs. The model’s input consists of a recent, continuous sub-sequence of the user’s lifelong history, along with a target item. For positive samples, the target items are the actual items users clicked next, and the model is expected to output “Yes.” For negative samples, the target items are randomly sampled, and the model should output “No.” In addition to these public datasets, we validated our performance on one of the world’s largest online advertising platforms. More datasets information and details such as training/validation/test splits are shown in Appendix B.1.
@@
We compare against a variety of recommendation models, including ETA (Chen et al., 2021), SDIM (Cao et al., 2022), DIN (Zhou et al., 2018), TWIN (Chang et al., 2023) and its variants, as well as TWIN-V2 (Si et al., 2024). As discussed in Sec.3.2, the target-aware representation by crossing $\bm{e}_{i}^{\text{Repr}}\odot\bm{v}_{t}^{\text{Repr}}$ significantly improves representation discriminability, so we include it in our baselines for fairness. TWIN-4E refers to the model introduced in Sec. 3.4, while TWIN (w/ proj.) refers to the model described in Sec. 2.3. TWIN (hard) represents a variant using “hard search” in the search stage, meaning it only retrieves behaviors with the same category as the target. TWIN (w/o TR) refers to the original TWIN model without target-aware representation, i.e., representing user history as $\bm{h}=\sum_{i}w_{i}\cdot\bm{e}_{i}$ instead of $\bm{h}=\sum_{i}w_{i}(\bm{e}_{i}\odot\bm{v}_{t})$.
@@ 4.2 Overall Performance
There are also some notable findings. TWIN outperforms TWIN (w/o TR) in most cases, proving that target-aware representation $\bm{e}_{i}^{\text{Repr}}\odot\bm{v}_{t}^{\text{Repr}}$ do help enhance discriminability (further evidence shown in Sec. 4.4). Our DARE model has an obvious advantage over TWIN-4E, confirming that the prior knowledge discussed in Sec. 3.4 is well-suited for the recommendation system. ETA and SDIM, which are based on TWIN and focus on accelerating the search stage at the expense of performance, understandably show lower AUC scores. TWIN-V2, a domain-specific method optimized for video recommendations, is less effective in our settings.
@@ 4.3 Attention Accuracy
Mutual information, which captures the shared information between two variables, is a powerful tool for understanding relationships in data. We calculate the mutual information between behaviors and the target as the ground truth correlation, following  (Zhou et al., 2024). The learned attention score reflects model’s measurement of the importance of each behavior. Therefore, we compare the attention distribution with mutual information in Fig. 9.
@@
The retrieval in the search stage relies entirely on attention scores. Thus, we further investigate the retrieval results on the test dataset, which provide a more intuitive reflection of attention quality. Behaviors with top-k mutual information are considered the ground truth for optimal retrieval, and we evaluate model performance using normalized discounted cumulative gain (NDCG) (Järvelin & Kekäläinen, 2002). The results, along with case studies, are presented in Fig. 10. We find that:
@@
  • TWIN is overly sensitive to temporal information. As discussed, TWIN tends to select recent behaviors regardless of their categories, against the ground truth, due to overestimated correlations between different categories, as shown in Fig. 10b and 10c.
  • TWIN with projection performs unstably. As shown in Sec. 2.3, the TWIN variant with projection has a highly dispersed attention logits distribution, with maximum logits often exceeding 10. This excessive “confidence” can be a double-edged sword. In some cases, like Fig. 10b, it performs exceptionally well; but in others, like Fig. 10c, it fails due to an “overconfident misjudgment” of an unimportant behavior.
Result 1. DARE succeeds in capturing the semantic-temporal correlation of behaviors and retaining correlated behaviors during the search stage, while other methods fail to do so.
@@
Figure 10: Retrieval in the search stage. (a) Our model can retrieve more correlated behaviors. (b-c) Two showcases where the x-axis is the categories of recent ten behaviors.
@@ 4.6 Online A/B Testing and Deployments
We apply our methods to one of the world’s largest online advertising platforms. In online advertising, users’ behaviors on ads are sparse, making the sequence length relatively shorter than the content recommendation scenario. To this end, we involve the user’s behavior sequence from our article and the micro-video recommendation scenario. Specifically, the user’s ad and content behaviors in the last two years are introduced. Before the search, the maximal length of the ads and content sequence is 4000 and 6000, respectively, while the average length is 170 and 1500, respectively. After searching using the proposed DARE, the sequence length is reduced to less than 500.
Regarding sequence features (side info), we choose the category ID, behavior type ID, scenario ID, and two target-aware temporal encodings, i.e., position relative to the target, and time interval relative to the target (with discretization). There are about 1.0 billion training samples per day. During the 5-day online A/B test in September 2024, the proposed DARE method achieves 0.57% cost, and 1.47% GMV (Gross Merchandize Value) lift over the production baseline of TWIN. This would lead to hundreds of millions of dollars in revenue lift per year.
4.7 Short-Sequence modeling
We have also tried our method in the short-sequence modeling, using the Amazon dataset (He & McAuley, 2016; McAuley et al., 2015) with the same setup as the state-of-the-art TIN model (Zhou et al., 2024). However, the performance improvement is marginal (TIN: 0.86291$\pm$0.0015 AUC vs. DARE: 0.86309$\pm$0.0004 AUC). This is likely because, unlike long-sequence modeling, short-sequence modeling lacks a search stage. As shown by Zhou et al. (2024), representation is more critical than attention in short-sequence settings, so the dominance of representation doesn’t significantly impact performance as long as all behaviors are preserved. In contrast, in long-sequence modeling, representation dominance affects retrieval during the search stage, causing some correlated behaviors to be filtered out and their representation lost.
@@ 5 Related Work
The attention mechanism, first introduced in Transformers (Vaswani et al., 2017), has proven highly effective and is widely used for correlation measurement. Transformers employ Q, K (attention projection), and V (representation projection) matrices to generate queries, keys, and values for each item. The scaled dot product of query and key serves as the correlation score, while the value serves as the representation. This structure is widely used in many domains, including natural language processing (Brown, 2020) and computer vision (Dosovitskiy, 2020). However, in recommendation systems, due to the interaction-collapse theory pointed out by Guo et al. (2023), the small embedding dimension would make linear projections completely lose effectiveness, as discussed in Sec. 2.3. Thus, proper adjustment is needed in this specific domain.
@@ A.1 Hyper-parameters and Model Details
Weight decay $10^{-6}$
The search stage retrieves 20 behaviors. Besides, we use the Adam optimizer. Layers of the Multi-layer Perceptron (MLP) are set as $200\times 80\times 2$, which is the same as Zhou et al. (2024).
@@ A.2 Baseline Implementation
  • DIN is a short-sequence modeling method, so we introduced the search stage and aligned it with long-sequence models. To be specific, the original DIN will aggregate all history behaviors based on a learned weight, so we let DIN also select top-K behaviors like other methods according to its learned weight. Note that the original DIN is not practical in long-sequence modeling, since aggregating such a long history is unacceptable due to the too high time complexity.
  • TWIN-V2 is specially designed for Kuaishou, a short video-sharing app, capturing video-specific features to make certain optimization, so we retained their core ideas and adjusted to our scenarios. We made some necessary adjustments, e.g., TWIN-V2 would first group the items based on the proportion a video is played, which does not have a corresponding feature in our datasets. So, we first group user history based on temporal information.
@@ B.1 Dataset Details
Some detailed information is shown in Table 2. We use Taobao (Zhu et al., 2018; 2019; Zhuo et al., 2020) and Tmall (Tianchi, 2018) datasets in our experiments. The proportion of active users in these two datasets is more than 60%, which is relatively more satisfying. Note that Taobao is relatively more complex, with more categories and more items, which is a higher challenge for model capacity.
@@
We sequentially number history behaviors from one (the most recent behavior) to T (the most ancient behavior) according to the time step. The test dataset contains predictions of the first behaviors, while the second behaviors are used as the validation dataset. For the training dataset, we use the $(3+5i,0\leq i\leq 18)$th behavior. Models would finish predicting the $j$th behavior based on $j-200$ to $j-1$ behaviors (padding if history length is not long enough). Only active users with behavior sequences longer than 210 will be reserved.

### v2

@@ 4.1 Setup
We use the publicly available Taobao (Zhu et al., 2018; 2019; Zhuo et al., 2020) and Tmall (Tianchi, 2018) datasets, which provide users’ behavior data over specific time periods on their platforms. Each dataset includes the items users clicked, represented by item IDs and their corresponding category IDs. Thus, a user’s history is modeled as a sequence of item and category IDs. The model’s input consists of a recent, continuous sub-sequence of the user’s lifelong history, along with a target item. For positive samples, the target items are the actual items users clicked next, and the model is expected to output “Yes.” For negative samples, the target items are randomly sampled, and the model should output “No.” In addition to these public datasets, we validated our performance on one of the world’s largest online advertising platforms. More details on datasets and training/validation/test splits are shown in Appendix B.
@@
We compare against a variety of recommendation models, including ETA (Chen et al., 2021), SDIM (Cao et al., 2022), DIN (Zhou et al., 2018), TWIN (Chang et al., 2023) and its variants, as well as TWIN-V2 (Si et al., 2024). As discussed in Sec.3.2, the target-aware representation by crossing $\bm{e}_{i}^{\text{Repr}}\odot\bm{v}_{t}^{\text{Repr}}$ significantly improves representation discriminability, so we include it in our baselines for fairness. TWIN-4E refers to the model introduced in Sec. 3.4, while TWIN (w/ proj.) refers to the model described in Sec. 2.3. TWIN (hard) represents a variant using “hard search” in the search stage, meaning it only retrieves behaviors from the same category as the target. TWIN (w/o TR) refers to the original TWIN model without target-aware representation, i.e., representing user history as $\bm{h}=\sum_{i}w_{i}\cdot\bm{e}_{i}$ instead of $\bm{h}=\sum_{i}w_{i}(\bm{e}_{i}\odot\bm{v}_{t})$.
@@ 4.2 Overall Performance
There are also some notable findings. TWIN outperforms TWIN (w/o TR) in most cases, proving that target-aware representation $\bm{e}_{i}^{\text{Repr}}\odot\bm{v}_{t}^{\text{Repr}}$ do help enhance discriminability (further evidence is shown in Sec. 4.4). Our DARE model has an obvious advantage over TWIN-4E, confirming that the prior knowledge discussed in Sec. 3.4 is well-suited for the recommendation system. ETA and SDIM, which are based on TWIN and focus on accelerating the search stage at the expense of performance, understandably show lower AUC scores. TWIN-V2, a domain-specific method optimized for video recommendations, is less effective in our settings.
@@ 4.3 Attention Accuracy
Mutual information, which captures the shared information between two variables, is a powerful tool for understanding relationships in data. We calculate the mutual information between behaviors and the target as the ground truth correlation, following  (Zhou et al., 2024). The learned attention score reflects the model’s measurement of the importance of each behavior. Therefore, we compare the attention distribution with mutual information in Fig. 8.
@@
The retrieval in the search stage relies entirely on attention scores. Thus, we further investigate the retrieval on the test dataset, which provides a more intuitive reflection of attention quality. Behaviors with top-k mutual information are considered the optimal retrieval, and we evaluate model performance using normalized discounted cumulative gain (NDCG) (Järvelin & Kekäläinen, 2002). The results, along with case studies, are presented in Fig. 9 (more examples in Appendix E.4). We find:
@@
  • TWIN is overly sensitive to temporal information. As discussed, TWIN tends to select recent behaviors regardless of their categories, against the ground truth, due to overestimated correlations between different categories, as shown in Fig. 9b and 9c.
  • Other methods perform unstably. For the other methods, they filter out some important behaviors and retrieve unrelated ones in many cases, which explains their bad performance.
Result 1. DARE succeeds in capturing the semantic-temporal correlation between behaviors and the target, retaining more correlated behaviors during the search stage.
@@
Figure 9: Retrieval in the search stage. (a) Our model can retrieve more correlated behaviors. (b-c) Two showcases where the x-axis is the categories of the recent ten behaviors.
@@ 4.6 Online A/B Testing and Deployments
We apply our methods to Tencent’s advertising platform. Since users’ behaviors on ads are sparse, which makes the sequence length relatively shorter than the content recommendation scenario, we involve the user’s behavior sequence from our article and the micro-video recommendation scenario. Specifically, the user’s ad and content behaviors in the last two years are introduced. Before the search, the maximal length of the ads and content sequence is 4000 and 6000, respectively, with 170 and 1500 on average. After searching with DARE, the sequence length is reduced to less than 500. Regarding sequence features (side info), we choose the category ID, behavior type ID, scenario ID, and two target-aware temporal encodings, i.e., position relative to the target, and time interval relative to the target (with discretization). There are about 1.0 billion training samples per day. During the 5-day online A/B test in September 2024, the proposed DARE method achieves 0.57% cost, and 1.47% GMV (Gross Merchandize Value) lift over the production baseline of TWIN. This would lead to hundreds of millions of dollars in revenue lift per year.
4.7 Supplementary Experiment Results in Appendix
Retrieval number in the search stage. DARE’s advantage is more obvious with less retrieval number, proving once again that DARE selects important behaviors more accurately (Appendix D.1).
Sequence length and short-sequence modeling. DARE can consistently benefit from longer sequences, while it delivers marginal advantages in short-sequence modeling (Appendix D.2).
GAUC and Logloss. Besides AUC, we also evaluate DARE and all the baselines under GAUC and Logloss. DARE shows consistent superiority, proving the solidity of our results (Appendix E.1).
@@ 5 Related Work
The attention mechanism, most well-known in Transformers (Vaswani et al., 2017), has proven highly effective and is widely used for correlation measurement. Transformers employ Q, K (attention projection), and V (representation projection) matrices to generate queries, keys, and values for each item. The scaled dot product of query and key serves as the correlation score, while the value serves as the representation. This structure is widely used in many domains, including natural language processing (Brown et al., 2020) and computer vision (Dosovitskiy et al., 2021). However, in recommendation systems, due to the interaction-collapse theory pointed out by Guo et al. (2024), the small embedding dimension would make linear projections completely lose effectiveness, as discussed in Sec. 2.3. Thus, proper adjustment is needed in this specific domain.
@@ 6 Conclusion
Reproducibility Statement
To ensure reproducibility, we provide the hyperparameters and baseline implementation details in Appendix A, along with dataset details in Appendix B. We have released the full code, including dataset processing, model training, and analysis experiments, at https://github.com/thuml/DARE.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (62021002), the BNRist Project, the Tencent Innovation Fund, and the National Engineering Research Center for Big Data Software.
@@ A.1 Hyper-parameters and Model Details
Weight decay 1e-6
Besides, we use the Adam optimizer. Layers of the Multi-layer Perceptron (MLP) are set as $200\times 80\times 2$, which is the same as Zhou et al. (2024).
@@ A.2 Baseline Implementation
  • DIN is primarily designed for short-sequence modeling, so we introduced the search stage and aligned it with long-sequence models. Specifically, while the original DIN aggregates all historical behaviors using a learned weight, our approach enables DIN to select the top-K most significant behaviors based on these weights, the same as other long-sequence modeling techniques. Note that the original DIN is impractical for long-sequence modeling, as aggregating such extensive history would result in prohibitively high time complexity.
  • TWIN-V2 is specifically designed for Kuaishou, a short video-sharing app, leveraging video-specific features to optimize performance in video recommendations. However, our experiments focus on a more general scenario where only item IDs and category IDs are available. Thus, we made some necessary adjustments while retaining the core ideas of TWIN-V2. e.g., TWIN-V2 would first group the videos based on the proportion a video is played, which does not have a corresponding feature in our datasets. Consequently, we grouped user history using temporal information instead. It’s understandable that outside its specific domain, TWIN-V2 cannot fully realize its potential.
@@ Appendix B Data Processing
Some detailed information is shown in Table 2. We use Taobao (Zhu et al., 2018; 2019; Zhuo et al., 2020) and Tmall (Tianchi, 2018) datasets in our experiments. The proportion of active users (Users with more than 50 behaviors) in these two datasets is more than 60%, which is relatively satisfying. Note that the Taobao dataset is more complex, with more categories and more items, which is a higher challenge for model capacity.
@@
We sequentially number history behaviors from one (the most recent behavior) to T (the most ancient behavior) according to the time step. The test dataset contains predictions of the first behaviors, while the second behaviors are used as the validation dataset. For the training dataset, we use the $(3+5i,0\leq i\leq 18)$th behavior. Models would finish predicting the $j$th behavior based on $j-200$ to $j-1$ behaviors (padding if history length is not long enough). Only users with behavior sequences longer than 210 will be reserved.